# Exploring a tentative link between FRBs and pulsars with broad energy distributions and applications for nearby FRB survey strategies

### v1

@@ 5 Conclusion
Before the discovery of a population of FRBs (Thornton et al., 2013) that supported the astrophysical origin of the “Lorimer burst” (Lorimer et al., 2007), pulsars (Manchester et al., 2005) were considered the primary sources of bright, short-duration ($\lesssim$ second) radio pulses, which require coherent emission (Pietka et al., 2015). While FRBs also necessitate coherent emission, their luminosities are more than ten orders of magnitude higher. Initially, the apparent one-off nature of FRBs led to favouring catastrophic scenarios (Zhang, 2023), distinguishing them from pulsars. However, the discovery of repeating FRBs (Spitler et al., 2016), and the investigation that some Galactic radio pulsing sources might be one-off (Keane, 2016; Zhang & Yang, 2024), have reduced this distinction. Further observations of fainter FRB bursts (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020; Nimmo et al., 2023; Zhang et al., 2024a) and brighter pulses from magnetars or pulsars (Kirsten et al., 2021; Good & Chime/Frb Collaboration, 2020; Bera & Chengalur, 2019) suggest no clear gap between the isotropic-equivalent energy releases of various radio pulses (Zhang et al., 2024a). Additionally, polarimetric similarities between FRBs and pulsars have been observed (Feng et al., 2022; Niu et al., 2024). These findings highlight the importance of exploring potential connections between FRBs and pulsars, at least for certain sources.
In this study, we have presented similarities in the high-energy distributions of two cosmological repeating FRBs (FRB 20201124A and FRB 20220912A), one nearby FRB 20200120E, and two pulsars (RRATs J1846--0257 and J1854+0306). All these sources exhibit power-law indices of $\alpha\gtrsim-1$, suggesting that their “bright pulses” might significantly contribute to the total radio pulse energy release. Furthermore, the brightest bursts from these sources fit well with a simple power-law model of $\alpha=-0.26\pm 0.05$. However, considering the limited sample size, we advise caution in interpreting this correlation as physical.
We encourage further observations to expand the sample of high-energy bursts from both FRB-like and pulsar-like sources. Current FRB surveys may be limited by selection effects, potentially missing FRBs with relative low-luminosity from the nearby GC systems (Zhang et al., 2024a). Survey strategies will need to vary for sources with different energy distributions. In this work, we have discussed detailed survey strategies for FAST, MeerKAT and Parkes cryoPAF for FRB in nearby GCs and suggested optimal survey targets. In principle, conducting combining observations for these targets using FAST for about $3$ hours and Parkes cryoPAF for $10-20$ hours could potentially lead to the discovery of new FRBs in nearby GCs, in practice. Additionally, a lack of detection after even a few hours of FAST observation can still provide strict constraints on flat $\alpha$ values.
@@ Acknowledgments
This research has been partially funded by the National SKA Program of China (2022SKA0130100), the National Natural Science Foundation of China (grant Nos. 12041306,12273113,12233002,12003028,12321003), the CAS Project for Young Scientists in Basic Research (Grant No. YSBR-063), the International Partnership Program of Chinese Academy of Sciences for Grand Challenges (114332KYSB20210018), the ACAMAR Postdoctoral Fellow, China Postdoctoral Science Foundation (grant No. 2020M681758), and the Natural Science Foundation of Jiangsu Province (grant Nos. BK20210998).

### v2

@@ 5 Conclusion
In this study, we have presented similarities in the high-energy distributions of two cosmological repeating FRBs (FRB 20201124A and FRB 20220912A), one nearby FRB 20200120E, and two pulsars (RRATs J1846-0257 and J1854+0306). All these sources exhibit power-law indices of $\alpha\gtrsim-1$, suggesting that their “bright pulses” might significantly contribute to the total radio pulse energy release. Furthermore, the brightest bursts from these sources could be fitted with a power-law model of $\alpha_{\rm Bri}=-0.26\pm 0.05$. Although we will not suggest interpreting this correlation as a physical relationship, the trend suggests that an extremely flat index is required to naturally derive high-luminosity radio bursts, such as FRBs, from low-luminosity bursts, such as those from pulsars. We encourage further observations to expand the sample of high-energy bursts from both FRB-like and pulsar-like sources.
As discussed in Section 4.2, current FRB surveys may miss FRBs with relatively low luminosity and narrow widths from nearby GC systems. Survey strategies will need to vary to search for sources with different energy distributions. In this work, we have discussed detailed survey strategies for FAST, MeerKAT and Parkes cryoPAF for FRB searches in nearby GCs and suggested optimal survey targets. In principle, conducting combined observations for these targets using FAST for about $3$ hours and Parkes cryoPAF for 102010-20 hours could potentially lead to the discovery of new FRBs in nearby GCs, in practice. Additionally, a lack of detection after even a few hours of FAST observation can still provide strict constraints on the combination of $\alpha$ values and intrinsic activity rates, or FRB rates.
@@ Acknowledgments
We thank the reviewer for the valuable and detailed comments, which have greatly improved the quality of this manuscript. This research has been partially funded by the International Partnership Program of Chinese Academy of Sciences for Grand Challenges (114332KYSB20210018), the National SKA Program of China (2022SKA0130100), the National Natural Science Foundation of China (grant Nos. 12041306,12273113,12233002,12003028,12321003), the CAS Project for Young Scientists in Basic Research (Grant No. YSBR-063), , and the ACAMAR Postdoctoral Fellow.
Supplementary Materials
Full details of all 1,832 Parkes cryoPAF pointings, which cover more than 7,000 clean GCs, are provided in the file “Parkes_PAF_pointing_above7000.txt”.